# Accessing Numerical Energy Hessians with Graph Neural Network Potentials and Their Application in Heterogeneous Catalysis

### v1

@@ 2.5 Complete Potential Energy Sampling
To facilitate the evaluation of the translational entropy, we created a suite of tools to perform potential energy sampling by the procedure described by Jørgensen et al9. The tool takes as input the adsorbate, slab, and an ASE 46 calculator object to perform constrained relaxations of the adsorbate on the slab. There is an ASE calculator object compatible with evaluating OC20 pretrained models, making their use facile. The constrained relaxations are performed on a grid with adjustable spacing. Results are cached so they may be accessed at a later time. Once the constrained relaxations are performed, a surrogate PES is constructed and used to evaluate the translational partition function and entropy by Equations 5 and 6, which was taken from the work by Jørgensen et al9. We also created classes similar to the ASE HarmonicThermo class, to allow for the rapid evaluation of the Gibbs free energy for a constructed PES and vibrational modes.
@@ 3.4 Complete potential energy sampling
For a subset of 17 systems which did not have a significant proportion of desorbed or errant systems, DFT single points were performed on the ML relaxed structures. A comparison of the relative DFT single point energies (relative to the minimum) and the relative ML energies can be seen in the lower, right of Figure 4. The majority of system have good energetic parity, but there are a few exceptions. For system B, there are a large number of points which are determined to be equal in energy by DFT, but ML erroneously finds the energies to have a distribution of values. For system A, which is shown in pink, the minimum energy by is erroneously small and far away from the distribution. This is the reason that the primary cluster of data is shifted away from the parity line but still correlated. The impact these artifacts have on the translational entropy component to the Gibbs free energy is shown in the bottom, center of Figure 4. The differences with system B have a minute impact on the overall correction, while the differences with system A have a large impact. Similar artifacts can be seen for the OC20Dense dataset21 for the GemNet-OC model18, and a figure supporting this has been included in the Supplementary Information. The dark blue points show the resultant correction when the erroneously small energies are kept, while the light blue points show the resultant correction when energies very far from the mean of the energy distribution are excluded. With the improvement brought about by enforcing energies to be in the distribution (within 2.5 standard deviations of the mean), the MAE between the DFT single points and the ML energies is just 0.016 eV at 300 K. This improvement is, however, modest for the system labeled A, and it is still a large outlier in the center Figure.
@@
Figure 4: Two examples of the complete potential energy sampling (CPES) technique for COH on an Pt-Sb-Zr alloy and NO3 on an Pd-Mn alloy. For each, the slab (center), potential energy surface (right), and the entropy profiles versus the temperature is shown for the harmonic approximation (HA) and CPES (left). The distribution of the translational component to the entropy correction evaluated at 300 K across 100 randomly sampled systems from validation (bottom, left). The parity between ML determined and DFT validated single point translational entropy Gibbs corrections (bottom, center). The parity between the relative adsorption energies used to construct the PES from ML and DFT single points (bottom, left).

### v2

@@ 2.5 Complete Potential Energy Sampling
To facilitate the evaluation of the translational entropy, we created a suite of tools to perform potential energy sampling by the procedure described by Jørgensen et al9. These tools are available on GitHub at https://github.com/jmusiel/gibby/tree/main. The tool takes as input the adsorbate, slab, and an ASE 46 calculator object to perform constrained relaxations of the adsorbate on the slab. There is an ASE calculator object compatible with evaluating OC20 pretrained models, making their use facile. The constrained relaxations are performed on a grid with adjustable spacing. Results are cached so they may be accessed at a later time. Once the constrained relaxations are performed, a surrogate PES is constructed and used to evaluate the translational partition function and entropy by Equations 5 and 6, which was taken from the work by Jørgensen et al9. We also created classes similar to the ASE HarmonicThermo class, to allow for the rapid evaluation of the Gibbs free energy for a constructed PES and vibrational modes.
@@ 3.4 Complete potential energy sampling
For a subset of 17 systems which did not have a significant proportion of desorbed or errant systems, DFT single points were performed on the ML relaxed structures. A comparison of the relative DFT single point energies (relative to the minimum) and the relative ML energies can be seen in the lower, right of Figure 4. The majority of systems have good energetic parity, but there are a few exceptions. For system B (orange), there are a large number of points which are determined to be equal in energy by DFT, but ML erroneously finds the energies to have a distribution of values. For system A, which is shown in pink, the minimum energy by is erroneously small and far away from the distribution. This is the reason that the primary cluster of data is shifted away from the parity line but still correlated. The impact these artifacts have on the translational entropy component to the Gibbs free energy is shown in the bottom, center of Figure 4. The differences with system B have a minute impact on the overall correction, while the differences with system A have a large impact. Similar artifacts can be seen for the OC20Dense dataset21 for the GemNet-OC model18, and a figure supporting this has been included in the Supplementary Information. There are two large outliers when comparing the translational entropy terms by ML and supplemented with DFT single points. Both of these were caused by the same issue described for system A: having a few ML energies that are far away from the energy distribution. The other outlier can not be seen the bottom, right of Figure 4 because the overall magnitude of its energy distribution is small so it is masked by the noise in other systems. In these cases, the translational entropy has been underestimated by ML. Still, the MAE in the translational entropy term by ML with reference to the DFT single points is just 0.023 eV at 300 K.
@@
Figure 4: Two examples of the complete potential energy sampling (CPES) technique for COH on an Pt-Sb-Zr alloy and NO3 on an Pd-Mn alloy. For each, the slab (center), potential energy surface (right), and the entropy profiles versus the temperature is shown for the harmonic approximation (HA) and CPES (left). The distribution of the translational component to the entropy correction evaluated at 300 K across 100 randomly sampled systems from validation (bottom, left). The parity between ML determined and DFT single point validated translational entropy Gibbs corrections (bottom, center). The parity between the relative adsorption energies used to construct the PES from ML and DFT single points (bottom, right). In the bottom, center and bottom, right figures two systems (A and B) have been highlighted so the impact of outliers in the relative energies (right) on the translational entropy term (center) can be seen.